# The Taylor-Wiles method for coherent cohomology, II

### v2

@@ 7. Proof of Theorem 24 (iii) and (iv)
The proof that the Galois representations attached to the boundary cohomology group $R\Gamma^{\partial}({\mathcal{F}}_{K}):=R\Gamma^{\partial}({\mathbb{S}}_{K}(G,X),{\mathcal{F}}_{K})$, or to its generalization $R\Gamma^{\partial}({\mathbb{S}}_{K}(G,X),{\mathcal{F}}_{K,\mbox{ю}})$, are reducible is based on a lengthy analysis of the toroidal boundary, to show that the Galois representations attached to any piece of the complex that computes the cohomology of the toroidal boundary breaks up as a sum of representations attached to the cohomology of smaller groups. This is a geometric argument and it is identical with or without the introduction of a $K$-type indicated by the subscript ю. Thus we will write down the proof, which is long enough as it is, in the case where the type data are trivial.
@@
There is a Shimura datum $(G_{h,R},X(R))$ attached to $R$, of the same kind as the pair $(G,X)$ with which we began. As noted in the Appendix, this is not quite the boundary Shimura datum that corresponds to $R$ in the theory of [53] or [31, 33]. In those references the group $G_{h,R}$ contains an additional split torus, here relegated to $G_{\ell,R}$, that is needed in order to define the correct mixed Hodge structure on the boundary cohomology. For the purposes of this paper this precision will not be necessary. The absence of this extra factor is compensated by the Tate twists in (66). [^5] Pink’s theory of mixed Shimura data $(G,X)$ allows for the possibility that $X$ is a union of a finite set of copies of hermitian symmetric domains in a flag variety attached to $G$. Taking this into account requires adjustments that are discussed in [33, §1.1.7], and that will be irrelevant for our purposes. We note for future reference that
@@
Since $\tilde{{\mathcal{W}}}$ is algebraic, it is defined over finite extension $E(\tilde{{\mathcal{W}}})$, and under our assumption admits a model over any ${\mathcal{O}}$-algebra $A$. For an open subgroup $K^{\prime}_{R}\subseteq K_{R}$, we have a natural finite covering map
@@ 7.1. Eisenstein classes
[33, Corollary 3.2.9] There is a spectral sequence abutting to the boundary cohomology in characteristic zero of the automorphic vector bundle ${\mathcal{F}}$:
@@
be the (unnormalized) partial Satake transform [50, §2.2.6]. Here “unnormalized" means the result of integration along the unipotent radical of $R$, but without multiplication by the modulus factor $\delta_{R}^{-1/2}$.
@@ 7.1.1. Proof of Claim 61
Let $\bar{r}_{i}^{d}(\nu^{\prime}_{\ell})=\bar{r}_{i}(\nu^{\prime}_{\ell})^{\vee}(1-m_{i}(R))$ (Tate twist). It then follows from identity (b) of Claim 60, as in the discussion following [50, Lemma 4.6], that there are integers $\mu_{0}$ and $\mu_{1}^{\pm},\mu_{2}^{\pm},\dots,\mu_{r(R)}^{\pm}$ such that $\nu^{\prime}_{R}$ and
@@ 7.2. Proof of Claim 58
of normal schemes. We assume ${\mathbb{S}}_{K,\Sigma}$ to be smooth and projective for convenience. The proof of [24, Lemma 2.6] applies in the mixed characteristic situation and implies that
@@
The first claim, as noted, follows as in the proof of [24, Lemma 2.6]. The first two spectral sequences then follow from Proposition 17 (iii), and the third from the definition of $R\Gamma^{\partial}$. ∎
@@ 7.2.1. Proof of Claim 58
(compare [50, Lemma 4.2]). For each $T$, the Hecke algebra $T_{T,K_{T}}$ acts on the $K_{T}$-invariants on the left-hand side compatibly with the action on $H^{\bullet}(R\Gamma^{\partial}_{\infty}({\mathcal{F}}_{K}))^{K_{T}}$ on the right hand side. Claim 58 now follows by combining (72) with (57).
@@ 8.1. The algebraic part
denote the inclusion of the corresponding locally closed stratum of the toroidal boundary, $\bar{Z}_{\sigma}$ its closure in ${\mathbb{S}}_{K}(G,X)_{\Sigma}$. Then we have a diagram as in [31, (1.2.5)], [45, (4.4),(4.5)] of smooth schemes over $Spec({\mathcal{O}})$:
@@
[45, Proposition 5.6] (i) There is a canonically determined locally free coherent sheaf ${\mathcal{F}}^{A}_{K}$ over $A_{F}$ and a canonical isomorphism
@@
The $P_{F}({{\mathbf{A}}_{f}^{p}})$-invariance of the filtration follows from the fact that the construction of automorphic vector bundles in [43, Definition 6.7] is a functor from locally free ${\mathcal{O}}$-representations of $P_{F}$ to $P_{F}({{\mathbf{A}}_{f}^{p}})$-equivariant vector bundles over $A_{F}$. The $P_{F}({{\mathbf{A}}_{f}^{p}})$-action is not mentioned explicitly in [43] but the construction is exactly analogous to that in [44, Proposition 8.1.4.1].
@@
The main reference for this fact is Theorem 2.15 of [43]. Since the notation $A_{F,\Sigma^{\prime}}$ is not quite justified there, Lan advises us to add a reference to Section 1.3.4 of [44]. Although the title of [44, §1.3] is “Algebraic compactifications in characteristic zero," the constructions in the relevant section are valid in mixed characteristic. ∎
@@
(ii) We first observe that each $R^{k}{\bar{\pi}_{1}}{\mathrm{gr}}^{j}({\mathcal{F}}^{A,\rm can}_{K})$ is a locally free sheaf on $\bar{M}_{F}$. Indeed, by (i), each ${\mathrm{gr}}^{j}({\mathcal{F}}^{A,\rm can}_{K})$ is a pullback via $\bar{\pi}_{1}$, so by the projection formula the claim reduces to the corresponding assertion for the structure sheaf of $A_{F,\Sigma^{\prime\prime}}$, where it follows from [43, Theorem 2.15]. The same result of Lan implies that our statement is true outside a divisor on the special fiber, and we know by the results of [31] that the statement is true on the generic fiber, and is true outside a divisor on the special fiber by results of Lan. Thus, as in the proof of Proposition 76, it is true globally, since $\bar{M}_{F}$ is a regular scheme. ∎
@@ 8.2. The topological part
Part (i) is proved as in [31, 33]. We compute the cohomology by the spectral sequence for the closed cover of $Z_{\Sigma}(R)$ by the maximal strata $\bar{Z}_{\sigma}$ as $\sigma$ runs over the simplices in $\mathfrak{N}_{\Sigma}(R)$. Proposition 76 allows us to apply the proof of [33, Proposition 3.1.1 (i)]. Part (ii) is a consequence Proposition 78. Part (iii) follows from the projection formula, in view of the isomorphism
@@
which is proved as [33, Proposition 3.4.1] in light of the results of this section and
@@ 8.2.3. The local system on $X(\Gamma^{red}_{\ell,R})$
Since the tensor category of automorphic vector bundles on $M_{F}$ is generated by the relative cotangent bundle of $A_{F}$ over $M_{F}$, this follows directly from the description of the latter in [44, (1.3.2.8)], given that the action of ${\Gamma}^{red}_{\ell,R}$ is linear on the factor denoted $X$ in that formula. Although the reference in [44] is asserted for the compactification in characteristic zero, the statements remain true in mixed characteristic as long as we are in a situation of good reduction, which is the case throughout this paper. ∎
@@ 8.3. Proof of Claim 60
for an automorphic vector bundle ${\mathcal{J}}_{K^{\prime}}$ on ${\mathbb{S}}_{K_{R}}(G_{h,R},X(R))$ and for some $a,b,c$. Part (a) of Claim 60 then follows by filtering ${\mathcal{J}}_{K^{\prime}}$ by irreducible subquotients. Part (b) is then standard. More precisely, it comes down to the following observation. Let $\ell\neq p$ be an unramified prime for $K$ and $\alpha$ be an irreducible spherical representation of $L_{R}({\mathbb{Q}}_{\ell})$ with coefficients in $k$; let $v_{R}\in\alpha$ be a non-zero spherical vector. Let $I(\alpha)$ denote the unnormalized induction of the pullback of $\alpha$ to $R({\mathbb{Q}}_{\ell})$ to $G({\mathbb{Q}}_{\ell})$, and let $v\in I(\alpha)$ be a (non-zero) spherical vector. Let $T(G,K_{\ell})$ and $T_{\ell}(L_{R},K_{R,\ell})$ denote the local spherical Hecke algebras of $G({\mathbb{Q}}_{\ell})$ and $L_{R}({\mathbb{Q}}_{\ell})$, respectively, relative to the indicated hyperspecial maximal compact subgroups. Then $T(G,K_{\ell})$ (resp. $T_{\ell}(L_{R},K_{R,\ell})$) acts on $v$ (resp. on $v_{R}$) through a character
@@
where $s_{R}$ is the unnormalized partial Satake transform, as above. See [50, Proposition 3.8] for the analogous case of topological cohomology, and see [22, Lemma 2.3.2] for the analogous global argument for holomorphic modular forms.
@@ A.2. Parametrization of adelic boundary components
The factorization (88) is a direct product. This implies that the groups denoted $\Delta_{1,R}$ and $\Delta_{0,R}$ that appear in the formula [33, (3.2.8)] are trivial. It also implies that the combinatorial structure of the toroidal boundary bears a simple relation to the topology of the Borel-Serre compactifications of the locally symmetric spaces attached to $G_{\ell,P}$, as $P$ varies.
@@
Good references for this are [21, §4] (though this only treats totally real fields the method is general), [54, §1.3.9], and [50], who consider torsion coefficients, as well as the unpublished book [20]. Fix once and for all a rational minimal parabolic $P_{0}\subset G$ with Levi decomposition $P_{0}=L_{0}\cdot U_{0}$. We consider open compact subgroups $K\subset G({{\mathbf{A}}_{f}})$ with the following properties:
@@ Acknowledgements
We thank Wushi Goldring, Vincent Pilloni, Jack Thorne, and Chandrashekhar Khare for answering our questions about aspects of their work that we consulted in the course of writing this paper, and for helping to bring us up to date on the relevant literature. In particular, we thank Pilloni for pointing to two proofs in his papers of commutativity of Hecke algebras acting on integral coherent cohomology. The authors are especially grateful to Kai-Wen Lan, who patiently explained where we could find specific references in his papers for the many results without which nothing in this paper would have been possible. Finally, we thank the anonymous referee for a careful reading of an earlier version of this paper.

### v3

@@ 7. Proof of Theorem 24 (iii) and (iv)
The proof that the Galois representations attached to the boundary cohomology group $R\Gamma^{\partial}({\mathcal{F}}_{K}):=R\Gamma^{\partial}({\mathbb{S}}_{K}(G,X),{\mathcal{F}}_{K})$, or to its generalization
$R\Gamma^{\partial}({\mathbb{S}}_{K}(G,X),{\mathcal{F}}_{K,\mbox{ю}}),$
are reducible is based on a lengthy analysis of the toroidal boundary, to show that the Galois representations attached to any piece of the complex that computes the cohomology of the toroidal boundary breaks up as a sum of representations attached to the cohomology of smaller groups. This is a geometric argument and it is identical with or without the introduction of a $K$-type indicated by the subscript ю. Thus we will write down the proof, which is long enough as it is, in the case where the type data are trivial.
@@
There is a Shimura datum $(G_{h,R},X(R))$ attached to $R$, of the same kind as the pair $(G,X)$ with which we began. As noted in the Appendix, this is not quite the boundary Shimura datum that corresponds to $R$ in the theory of [55] or [32, 34]. In those references the group $G_{h,R}$ contains an additional split torus, here relegated to $G_{\ell,R}$, that is needed in order to define the correct mixed Hodge structure on the boundary cohomology. For the purposes of this paper this precision will not be necessary. The absence of this extra factor is compensated by the Tate twists in (66). [^5] Pink’s theory of mixed Shimura data $(G,X)$ allows for the possibility that $X$ is a union of a finite set of copies of hermitian symmetric domains in a flag variety attached to $G$. Taking this into account requires adjustments that are discussed in [34, §1.1.7], and that will be irrelevant for our purposes. We note for future reference that
@@
Since $\tilde{{\mathcal{W}}}$ is algebraic, it is defined over a finite extension $E(\tilde{{\mathcal{W}}})$, and under our assumption admits a model over any ${\mathcal{O}}$-algebra $A$. For an open subgroup $K^{\prime}_{R}\subseteq K_{R}$, we have a natural finite covering map
@@ 7.1. Eisenstein classes
[34, Corollary 3.2.9] There is a spectral sequence abutting to the boundary cohomology in characteristic zero of the automorphic vector bundle ${\mathcal{F}}$:
@@
be the (unnormalized) partial Satake transform [52, §2.2.6]. Here “unnormalized" means the result of integration along the unipotent radical of $R$, but without multiplication by the modulus factor $\delta_{R}^{-1/2}$.
@@ 7.1.1. Proof of Claim 61
Let $\bar{r}_{i}^{d}(\nu^{\prime}_{\ell})=\bar{r}_{i}(\nu^{\prime}_{\ell})^{\vee}(1-m_{i}(R))$ (Tate twist). It then follows from identity (b) of Claim 60, as in the discussion following [52, Lemma 4.6], that there are integers $\mu_{0}$ and $\mu_{1}^{\pm},\mu_{2}^{\pm},\dots,\mu_{r(R)}^{\pm}$ such that $\nu^{\prime}_{R}$ and
@@ 7.2. Proof of Claim 58
of normal schemes. We assume ${\mathbb{S}}_{K,\Sigma}$ to be smooth and projective for convenience. The proof of [25, Lemma 2.6] applies in the mixed characteristic situation and implies that
@@
The first claim, as noted, follows as in the proof of [25, Lemma 2.6]. The first two spectral sequences then follow from Proposition 17 (iii), and the third from the definition of $R\Gamma^{\partial}$. ∎
@@ 7.2.1. Proof of Claim 58
(compare [52, Lemma 4.2]). For each $T$, the Hecke algebra $T_{T,K_{T}}$ acts on the $K_{T}$-invariants on the left-hand side compatibly with the action on $H^{\bullet}(R\Gamma^{\partial}_{\infty}({\mathcal{F}}_{K}))^{K_{T}}$ on the right hand side. Claim 58 now follows by combining (72) with (57).
@@ 8.1. The algebraic part
denote the inclusion of the corresponding locally closed stratum of the toroidal boundary, $\bar{Z}_{\sigma}$ its closure in ${\mathbb{S}}_{K}(G,X)_{\Sigma}$. Then we have a diagram as in [32, (1.2.5)], [47, (4.4),(4.5)] of smooth schemes over $Spec({\mathcal{O}})$:
@@
[47, Proposition 5.6] (i) There is a canonically determined locally free coherent sheaf ${\mathcal{F}}^{A}_{K}$ over $A_{F}$ and a canonical isomorphism
@@
The $P_{F}({{\mathbf{A}}_{f}^{p}})$-invariance of the filtration follows from the fact that the construction of automorphic vector bundles in [45, Definition 6.7] is a functor from locally free ${\mathcal{O}}$-representations of $P_{F}$ to $P_{F}({{\mathbf{A}}_{f}^{p}})$-equivariant vector bundles over $A_{F}$. The $P_{F}({{\mathbf{A}}_{f}^{p}})$-action is not mentioned explicitly in [45] but the construction is exactly analogous to that in [46, Proposition 8.1.4.1].
@@
The main reference for this fact is Theorem 2.15 of [45]. Since the notation $A_{F,\Sigma^{\prime}}$ is not quite justified there, Lan advises us to add a reference to Section 1.3.4 of [46]. Although the title of [46, §1.3] is “Algebraic compactifications in characteristic zero," the constructions in the relevant section are valid in mixed characteristic. ∎
@@
(ii) We first observe that each $R^{k}{\bar{\pi}_{1}}{\mathrm{gr}}^{j}({\mathcal{F}}^{A,\rm can}_{K})$ is a locally free sheaf on $\bar{M}_{F}$. Indeed, by (i), each ${\mathrm{gr}}^{j}({\mathcal{F}}^{A,\rm can}_{K})$ is a pullback via $\bar{\pi}_{1}$, so by the projection formula the claim reduces to the corresponding assertion for the structure sheaf of $A_{F,\Sigma^{\prime\prime}}$, where it follows from [45, Theorem 2.15]. The same result of Lan implies that our statement is true outside a divisor on the special fiber, and we know by the results of [32] that the statement is true on the generic fiber, and is true outside a divisor on the special fiber by results of Lan. Thus, as in the proof of Proposition 76, it is true globally, since $\bar{M}_{F}$ is a regular scheme. ∎
@@ 8.2. The topological part
Part (i) is proved as in [32, 34]. We compute the cohomology by the spectral sequence for the closed cover of $Z_{\Sigma}(R)$ by the maximal strata $\bar{Z}_{\sigma}$ as $\sigma$ runs over the simplices in $\mathfrak{N}_{\Sigma}(R)$. Proposition 76 allows us to apply the proof of [34, Proposition 3.1.1 (i)]. Part (ii) is a consequence Proposition 78. Part (iii) follows from the projection formula, in view of the isomorphism
@@
which is proved as [34, Proposition 3.4.1] in light of the results of this section and
@@ 8.2.3. The local system on $X(\Gamma^{red}_{\ell,R})$
Since the tensor category of automorphic vector bundles on $M_{F}$ is generated by the relative cotangent bundle of $A_{F}$ over $M_{F}$, this follows directly from the description of the latter in [46, (1.3.2.8)], given that the action of ${\Gamma}^{red}_{\ell,R}$ is linear on the factor denoted $X$ in that formula. Although the reference in [46] is asserted for the compactification in characteristic zero, the statements remain true in mixed characteristic as long as we are in a situation of good reduction, which is the case throughout this paper. ∎
@@ 8.3. Proof of Claim 60
for an automorphic vector bundle ${\mathcal{J}}_{K^{\prime}}$ on ${\mathbb{S}}_{K_{R}}(G_{h,R},X(R))$ and for some $a,b,c$. Part (a) of Claim 60 then follows by filtering ${\mathcal{J}}_{K^{\prime}}$ by irreducible subquotients, one of which will be the ${\mathcal{F}}^{R}$ of the statement of Claim 60. Part (b) is then standard. More precisely, it comes down to the following observation. Let $\ell\neq p$ be an unramified prime for $K$ and $\alpha$ be an irreducible spherical representation of $L_{R}({\mathbb{Q}}_{\ell})$ with coefficients in $k$; let $v_{R}\in\alpha$ be a non-zero spherical vector. Let $I(\alpha)$ denote the unnormalized induction of the pullback of $\alpha$ to $R({\mathbb{Q}}_{\ell})$ to $G({\mathbb{Q}}_{\ell})$, and let $v\in I(\alpha)$ be a (non-zero) spherical vector. Let $T(G,K_{\ell})$ and $T_{\ell}(L_{R},K_{R,\ell})$ denote the local spherical Hecke algebras of $G({\mathbb{Q}}_{\ell})$ and $L_{R}({\mathbb{Q}}_{\ell})$, respectively, relative to the indicated hyperspecial maximal compact subgroups. Then $T(G,K_{\ell})$ (resp. $T_{\ell}(L_{R},K_{R,\ell})$) acts on $v$ (resp. on $v_{R}$) through a character
@@
where $s_{R}$ is the unnormalized partial Satake transform, as above. See [52, Proposition 3.8] for the analogous case of topological cohomology, and see [23, Lemma 2.3.2] for the analogous global argument for holomorphic modular forms.
@@ A.2. Parametrization of adelic boundary components
The factorization (88) is a direct product. This implies that the groups denoted $\Delta_{1,R}$ and $\Delta_{0,R}$ that appear in the formula [34, (3.2.8)] are trivial. It also implies that the combinatorial structure of the toroidal boundary bears a simple relation to the topology of the Borel-Serre compactifications of the locally symmetric spaces attached to $G_{\ell,P}$, as $P$ varies.
@@
Good references for this are [22, §4] (though this only treats totally real fields the method is general), [56, §1.3.9], and [52], who consider torsion coefficients, as well as the unpublished book [21]. Fix once and for all a rational minimal parabolic $P_{0}\subset G$ with Levi decomposition $P_{0}=L_{0}\cdot U_{0}$. We consider open compact subgroups $K\subset G({{\mathbf{A}}_{f}})$ with the following properties:
@@ Acknowledgements
We thank Wushi Goldring, Vincent Pilloni, Jack Thorne, Najmuddin Fakhruddin, and Chandrashekhar Khare for answering our questions about aspects of their work that we consulted in the course of writing this paper, and for helping to bring us up to date on the relevant literature. In particular, we thank Pilloni for pointing to two proofs in his papers of commutativity of Hecke algebras acting on integral coherent cohomology. The authors are especially grateful to Kai-Wen Lan, who patiently explained where we could find specific references in his papers for the many results without which nothing in this paper would have been possible. Finally, we thank the anonymous referee for a careful reading of an earlier version of this paper.